Two cars are traveling along a straight road. Car (A) maintains a constant speed of 70.0 km/h; car (B) maintains a constant speed of 106 km/h. At t = 0, car (B) is 36.0 km behind car A. How far will car (A) travel from t = 0 before it is overtaken by car (B)?
Answer #1

speed of car A =VA = 70 km/h

speed of car B = VB = 106 km/h

distance travelled by car A = SA = 70 * t

distance travelled by car B = SB = 106 *t

SB-SA = 36 km

(106 - 70) t = 36

t = 1 hour

SA = 70 * 1 = 70 Km = 70,000m

at